### 01 Mar 2024 v2

# Collapse of wave functions in Schrödinger’s wave mechanics

Abstract
We show that inelastic scattering leads to a collapse of the wave function within standard evolution through the Schrödinger equation, whereas elastic scattering will not collapse the wave function. Specifically, we find that the initial width of the emerging wave function in inelastic scattering is primarily determined by the size of the participating scattering center, but not by the width of the incoming wave function. This implies that dynamical collapse of the wave function through inelastic scattering, together with energy quantization in bound quantum systems, can explain the emergence of particle-like signals without the need to invoke the Born rule.
Wave function collapse, Born rule
The Born interpretation of the wave function $\psi(\bm{x},t)$ of a nonrelativistic particle asserts that the probability to find the particle in a volume $V$ at time $t$ is
$P_{V}(t)=\int_{V}\!d^{3}\bm{x}\,|\psi(\bm{x},t)|^{2}.$ (1)
Furthermore, the spatial extension of the signal from an individual particle will be determined by detector resolution, but not by the width of the probability density $|\psi(\bm{x},t)|^{2}$. This agrees with observations in single-particle diffraction experiments [1, 2, 3]. Interference effects generate macroscopic separations between different maxima of $|\psi(\bm{x},t)|^{2}$ and sampling the signals from many particles produces the interference patterns. However, every single particle signal is still pointlike in the sense of being determined by detector resolution.
These observations encode the “measurement problem” of quantum mechanics in a nutshell. The continuous evolution of the wave function according to the Schrödinger equation can produce mesoscopic or macroscopic distributions $|\psi(\bm{x},t)|^{2}$, and the confirmation of interference patterns from summation of many single-particle signals confirms that evolution of the single-particle wave function probes all the relevant scattering centers or paths on the way to the detector, and yet the final individual single-particle signal in the detector is pointlike and may involve interaction with only a single atom-scale scattering center (e.g. on a fluorescent screen), or a sequence of pointlike interactions with consecutive single scattering centers (e.g. in a cloud chamber or bubble chamber), thus producing a “classical” particle track.
Motion from an electron gun to an electron detector produces a wave function that corresponds to a coherent superposition of scattering amplitudes from several scattering centers (in order $e^{2}$ of the sum of amplitudes, i.e. the particle did not interact with one scattering center after the other), but the act of detection seems to break this coherent evolution and suddenly the particle interacts with only one scattering center in order $e^{2}$ of the final scattering amplitude. The very act of “observing” or “recording” the particle with a detector seems to break the continuous wave-like evolution of the wave function, although the particle had simultaneously probed several scattering centers over many atomic scales, or even over macroscopic distances, before the detection.
This led to the speculations about a special role of observers in the old Copenhagen interpretation of quantum mechancis: The combination of scattering and observation does not register any pre-existing orbit of the particle but creates that orbit [4, 5, 6, 7, 8]. Heisenberg took a softer stance later in life and argued that localization of the particle during the measurement process may also be a consequence of complexity of interactions of the particle during the detection, see Heisenberg’s remarks in [9], in particular p. 23.
The indeterminacy of particle properties is well documented in spin experiments and widely accepted textbook knowledge in quantum mechanics. However, any special role of detection (or observation or measurement or recording) of a particle in localizing the particle, due to the presence or the actions of a conscious observer, appears far fetched and did not become generally accepted in the physics community.
The assumption that localization happens due to complexity of interactions during the measurement process appears much more agreeable and contributed to the motivation for the decoherence program, which made important contributions to our understanding of the quantum-to-classical transition and the emergence of classically preferred states [10, 11, 12, 13, 14]. However, complexity of interactions cannot solve the measurement problem as such. The interaction of an electron with the fluorescent screen in a low energy electron diffraction (LEED) experiment is no more complex than its coherent interaction with the surface atoms in the sample. Both kinds of interactions are described in order $e^{2}$ through first-order scattering amplitudes of an electron interacting with a medium, and there is no physical rationale for arguing that scattering off the fluorescent screen should be more complex than scattering off the sample. Indeed, the designation “measurement problem” is a misnomer that reflects the fact that particle-like behavior was perceived to be linked to observations in the early development of quantum mechanics. However, this is clearly not the case. Once we trigger the electron gun in an LEED device, the electrons will scatter coherently off the sample and produce pointlike flashes on the screen, even when we do not watch nor turn on a CCD camera for recording.
It is tempting to infer from observation that the apparent break in wave function evolution is a consequence of different behavior of the wave function under elastic and inelastic scattering. Generation of interference effects from simultaneous contributions of several scattering centers or different paths requires elastic scattering according to the Laue conditions, while registration in a detector requires energy deposition through inelastic scattering. Inelastic scattering in a medium can generate pointlike signals, e.g. on a scintillation screen, or tracks in a liquid medium, e.g. in cloud chambers or bubble chambers.
Particle detection through inelastic scattering implies that we deliberately generate situations where inelastic scattering is likely to occur. However, the premise of science to identify laws of nature that desribe the universe independently from any conscious measurement or observation, implies that a particle will behave in exactly the same way irrespective of whether a particular medium is used for particle detection or not. We would posit that most scientists believe that a charged particle in a supersaturated vapor will generate a track in every corner of the universe, irrespective of how far away from the next life form or civilization. The question is then not whether observation or measurement changes smooth evolution according to the Schrödinger equation, but whether the Schrödinger equation itself can explain the emergence of particle tracks.
Mott [15] and Heisenberg [7] had pointed out that forward domination of Coulomb scattering can naturally extend a particle track where a track has been started. However, as Schonfeld correctly noted, the key question from the point of view of the measurement problem is the formation of the beginning of the track, the first droplet [16]. Based on properties of polarized molecules in supersaturated vapors, Schonfeld argues that Schrödinger evolution of the wave function should be able to explain the emergence of the track through a “consumption” or “draining” of the initial spherical wave through the scattered wave function.
However, we would submit that every medium that can generate repeated inelastic particle scattering, should generate a particle track, and this should not critically depend on properties of the scattering centers. We therefore model the medium as a set of scattering centers at locations $\bm{a}_{I}$, where $I$ enumerates the scattering centers. Furthermore, we assume fixed location of the scattering centers, as in a solid matrix. We are not aiming for realistic modeling of liquid-medium detectors. We are rather interested in a proof-of-principle investigation whether inelastic scattering can trigger wave function collapse within the scope of Schrödinger’s wave equation.
The general form of the model Hamiltonian for the particle-plus-medium system is then
$H$ $=$ $H_{0}+V$ (2)
$=$ $\frac{\bm{p}^{2}}{2m}+\sum_{I}\left(\frac{\bm{P}_{I}^{2}}{2M}+U(\bm{y}_{I}-\bm{a}_{I})\right)+\sum_{I}V(\bm{x}-\bm{y}_{I}),$
where $\bm{p}$ and $m$ are momentum and mass of the particle, and $\bm{P}_{I}$ and $M$ refer to momenta and mass of the scattering centers. The internal potentials $U(\bm{y}_{I}-\bm{a}_{I})$ generate the spectrum of the scattering centers and $V(\bm{x}-\bm{y}_{I})$ is the short-range scattering potential of the particle with the $I$-th scattering center. This would generically be a screened Coulomb potential in realistic models.
We also assume that the scattering centers before scattering are in their ground state with wave function $\phi_{0}(\bm{y}_{I}-\bm{a}_{I})$. For example, inelastic scattering of low-energy $\alpha$ particles (with energies of order of a few 100 keV to a few MeV) in a solid material primarily excites vibrations of the lattice of nuclei, because the electronic form factors for momentum transfers $\Delta p=\hbar\Delta k\gtrsim 100\,\mathrm{keV}/c$ satisfy $F(\Delta k)\ll 1$. Therefore, assuming harmonic oscillators of frequency $\omega$, $U(\bm{y}_{I}-\bm{a}_{I})=M\omega^{2}(\bm{y}_{I}-\bm{a}_{I})^{2}/2$, is not completely unrealistic as a first approximation for the scattering centers in this case.
The Hamiltonian would evolve an initial state $|\Psi(t^{\prime})\rangle=\exp(-\,\mathrm{i}H_{0}t^{\prime}/\hbar)|\Psi_{i}\rangle$ into the state $|\Psi(t)\rangle=\exp(-\,\mathrm{i}H_{0}t/\hbar)|\Psi_{f}\rangle$. Here both the initial and the final state of the particle-plus-medium system are expressed in terms of fiducial states $|\Psi_{i}\rangle$ and $|\Psi_{f}\rangle$ at time 00, where the fiducial states are related to the actual system state at times $t^{\prime}$ and $t$ through free time evolution, i.e. we describe the state before and after inelastic scattering through asymptotic free states. This complies with the assumption of finite range of the scattering potential $V$.
Canonical time evolution of $|\Psi(t)\rangle$ with the Schrödinger equation implies
$|\Psi_{f}\rangle$ $=$ $\exp(\mathrm{i}H_{0}t/\hbar)\exp[-\,\mathrm{i}H(t-t^{\prime})/\hbar]\exp(-\,\mathrm{i}H_{0}t^{\prime}/\hbar)|\Psi_{i}\rangle$ (3)
[…]
which is equivalent to the scattering matrix with the Dirac picture Hamiltonian
$V_{D}(t)=\exp(\mathrm{i}H_{0}t/\hbar)V\exp(-\,\mathrm{i}H_{0}t/\hbar).$ (4)
The connection with the scattering matrix
[…]
becomes explicit if we expand $|\Psi_{f}\rangle$ and $|\Psi_{i}\rangle$ in terms of eigenstates $|\Psi^{(0)}_{n}\rangle$ of $H_{0}$,
$|\Psi_{f}\rangle=\sum_{n,m}|\Psi^{(0)}_{n}\rangle S_{nm}(t,t^{\prime})\langle\Psi^{(0)}_{m}|\Psi_{i}\rangle.$ (6)
The scattering matrix contains all possible scattering channels of the particle-plus-medium system, and we focus on first-order scattering off the scattering potential $V$. First-order scattering already includes the possibility of elastic scattering from all the scattering centers. This maps the particle momentum $\bm{p}_{i}\to\bm{p}_{f}$ with $|\bm{p}_{f}|=|\bm{p}_{i}|$, and all coherently illuminated scattering centers contribute to the amplitude that takes $\bm{p}_{i}$ to $\bm{p}_{f}$. Inelastic scattering, on the other hand, leaves an imprint on the medium in exciting a scattering center in first order. The magnitudes of the scattering matrix elements determine which transitions are more likely to happen. Surfaces tested in LEED devices yield electron scattering matrices that are dominated by elastic terms, while any of the subdominant inelastic terms decrease the signal-to-noise ratio. Electron detection devices, on the other hand, should yield scattering matrices that are dominated by inelastic terms to detect electrons through energy transfer to individual scattering centers.
We assume that inelastic scattering excites the scattering center at location $\bm{a}_{I}$ from the ground state with wave function $\phi_{0}(\bm{y}_{I}-\bm{a}_{I})$ and energy $E_{0}$ into an excited state with wave function $\phi_{1}(\bm{y}_{I}-\bm{a}_{I})$ and energy $E_{1}$. The other scattering centers do not contribute in leading order of the scattering potential, and we can represent the state as a two-particle wave function $\langle\bm{x},\bm{y}_{I}|\Psi(t)\rangle$. The short range of the scattering potential then implies that we can write the two-particle state before and after the scattering in the form
$|\Psi(t^{\prime})\rangle$ $=$ $\exp(-\,\mathrm{i}H_{0}t^{\prime}/\hbar)|\psi_{i}\rangle\otimes|\phi_{0}\rangle,$ (7)
$|\Psi(t)\rangle$ $=$ $\exp(-\,\mathrm{i}H_{0}t/\hbar)|\psi_{f}\rangle\otimes|\phi_{f}\rangle.$ (8)
Free time evolution generates dispersion of wave packets with a time scale $\tau=2m\Delta x^{2}/\hbar$. However, the free dispersion time scale is much larger than particle travel times in lab experiments with electrons from millimetre aperture electron guns or neutrons from centimetre aperture neutron pipes, see, e.g. [17]. Our condition for emergence of a pointlike signal from wave function collapse is therefore that the width of $\langle\bm{x}|\psi_{f}\rangle$ is primarily determined by the size of the scattering centers, but not by the width of $\langle\bm{x}|\psi_{i}\rangle$.
Projection of the time evolution (3) into the single-particle sector yields in leading order of the scattering potential
[…]
×$\times\phi^{+}_{1}(\bm{y}-\bm{a}_{I})\phi_{0}(\bm{y}-\bm{a}_{I})\exp\!\left(-\,\mathrm{i}\frac{\bm{p}^{2}\tau}{2m\hbar}\right)|\psi_{i}\rangle,$
where $\bm{p}$ and $\bm{x}$ are the operators for the scattered particle. The transition frequency $\omega_{10}$ corresponds to the energy transfer from the particle to the scattering center,
$\omega_{10}=(E_{1}-E_{0})/\hbar.$ (10)
Evaluation of Eq. (9) in the usual limits $t\to\infty$, $t^{\prime}\to-\,\infty$ (which practically only means that $t-t^{\prime}$ should be large compared to the travel time of the particle through the range of the scattering potential), yields the outgoing wave packet $\psi_{f}(\bm{x})$ in terms of the incoming wave packet $\psi_{i}(\bm{x})$, both taken at fiducial time 00,
$\psi_{f}(\bm{x})$ $=$ $\frac{m}{(2\pi)^{3}\mathrm{i}\hbar^{2}}\int\!d^{3}\bm{y}\int\!d^{3}\bm{z}\int\!d^{3}\bm{x}^{\prime}\int\!d^{3}\bm{k}\,\exp[\mathrm{i}\bm{k}\cdot(\bm{x}-\bm{z})]$ (11)
×$\times\frac{\sin\!\left(\sqrt{k^{2}+(2m\omega_{10}/\hbar)}|\bm{z}-\bm{x}^{\prime}|\right)}{\pi|\bm{z}-\bm{x}^{\prime}|}\,V(\bm{z}-\bm{y})$
×$\times\phi^{+}_{1}(\bm{y}-\bm{a}_{I})\phi_{0}(\bm{y}-\bm{a}_{I})\psi_{i}(\bm{x}^{\prime}).$
The wave packet $\psi_{f}(\bm{x})$ is not yet normalized because it emerged from the unitary time evolution (3) through projection onto the excited state $\phi_{1}(\bm{y}-\bm{a}_{I})$ of the scattering center. This yields only a conditional probability for this particular inelastic scattering process relative to the other elastic and inelastic scattering channels that are possible in (3). However, we can infer interesting information on the width of inelastically scattered wave packets from Eq. (11).
The Dirichlet kernel $\sin\!\left(\kappa|\bm{z}-\bm{x}^{\prime}|\right)/(\pi|\bm{z}-\bm{x}^{\prime}|)$ in Eq. (11) provides an Ångström-scale approximation to a $\delta$-function already for $k=0$ and for particle masses $m\geq m_{e}=511\,\mathrm{keV}/c^{2}$ if the excitation energy satisfies $\hbar\omega_{10}\gtrsim 1\,\mathrm{eV}$. Furthermore, the factor $\phi^{+}_{1}(\bm{y}-\bm{a}_{I})\phi_{0}(\bm{y}-\bm{a}_{I})$ will provide atom-scale resolution around the location $\bm{a}_{I}$ of the scattering center, and screened Coulomb potentials have Ångström-scale range in dense materials. All this indicates that the emerging wave packet from inelastic scattering should be located at the scattering center and have an initial width that is determined by the width of the scattering center, in perfect agreement with the operational definition of a particle-like signal. Indeed, approximating the Ångström-scale factors in Eq. (11) with $\delta$-functions yields
$\psi_{f}(\bm{x})\sim\phi^{+}_{1}(\bm{x}-\bm{a}_{I})\phi_{0}(\bm{x}-\bm{a}_{I})\psi_{i}(\bm{x}).$ (12)
The atomic transition factor $\phi^{+}_{1}(\bm{x}-\bm{a}_{I})\phi_{0}(\bm{x}-\bm{a}_{I})$ will cut a piece of atom-scale resolution out of the wider wave function of the incoming particle, and this piece will be anchored around the scattering center. “Pointlike” position detection of subatomic particles per se is therefore not a miracle in wave mechanics.
We note that this derivation of wave function collapse works only for inelastic scattering. The scattered wave function for elastic scattering will always be a superposition of the 0th order term $\psi^{(0)}_{f}(\bm{x})\sim\psi_{i}(\bm{x})$ and a first order term $\psi^{(1)}_{f}(\bm{x})$. The 0th order term implies that the elastically scattered wave packet should at least inherit the width of the incoming wave packet. Furthermore, the Dirichlet kernel for elastic scattering contains no energy transfer term $2m\Delta E/\hbar^{2}$ in the argument, and this implies that the Dirichlet kernel will not provide an atom-scale approximation to a $\delta$-function for small particle momenta. The scattering matrix tells us that wave function collapse will happen in inelastic scattering, but not in elastic scattering.
However, wave function collapse in inelastic scattering by itself may not absolve us of the measurement problem in wave mechanics. The catch is that the mathematical evolution of the wave function through inelastic scattering might also allow for the superposition of two inelastic scattering events in locations $\bm{a}_{1}$ and $\bm{a}_{2}$. This would yield the superposition of a pointlike particle signal in location $\bm{a}_{1}$ due to excitation of the scattering center $\bm{a}_{1}$, with another pointlike particle signal in location $\bm{a}_{2}$ due to excitation of the scattering center $\bm{a}_{2}$. In the absence of any additional restrictions on permissible final states, coherent superposition of two inelastic scattering events appears to correspond to an allowed final state of the scattering matrix. With respect to the detector, this would generate a superposition of states proportional to $|\phi_{1,\bm{a}_{1}}\rangle|\phi_{0,\bm{a}_{2}}\rangle+\exp(\mathrm{i}\beta)|\phi_{0,\bm{a}_{1}}\rangle|\phi_{1,\bm{a}_{2}}\rangle$ (up to normalization, and with an arbitrary relative phase $\beta$). Without the Born rule, the Schrödinger equation might allow for the formation of more than one particle track, or more than one point lit on a fluorescent screen.
This is not a problem in epistemic interpretations of the wave function [18, 19, 20, 21, 22, 23] because the Born rule is understood as a necessary prescription for translating an epistemic wave function into a position signal.
It is also not a problem in Bohmian mechanics [24, 25, 26, 27] because the inclusion of a quantum potential and the addition of a guiding equation make the Born rule redundant.
Notwithstanding, wave function collapse in inelastic scattering opens up a third line of attack on the measurement problem. We may be able to exclude the simultaneous excitation of different scattering centers through the observation that low-energy excitations in isolated bound quantum systems can only yield energy eigenstates. This excludes states of the form [^1] We assume that we can neglect interactions between different scattering centers. $|\phi_{1,\bm{a}_{1}}\rangle|\phi_{0,\bm{a}_{2}}\rangle+\exp(\mathrm{i}\beta)|\phi_{0,\bm{a}_{1}}\rangle|\phi_{1,\bm{a}_{2}}\rangle$ as final detector states and explains the emergence of a single pointlike position signal without invoking the Born rule.
Acknowledgements.
We acknowledge support from the Natural Sciences and Engineering Research Council of Canada.